# Geometry-Aware Approaches for Balancing Performance and Theoretical Guarantees in Linear Bandits

### v1

@@ 4.1 An Instance-Dependent Regret Bound for POFUL
In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$. The following proposition bounds the instantaneous regret of POFUL and we defer the proof to Appendix D.2.
@@
On the RHS, since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\left\|\widetilde{x}_{t}\right\|_{V_{t}^{-1}}^{2}$ to get an upperbound of the regret. To fix the problem due to this discrepancy, the key point is connecting $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ in terms of the $V_{t}^{-1}$ norm.
@@
For each $t\geq 1$, let $\widetilde{x}_{t}$ and $x_{t}^{\star}$ respectively denote the action chosen by POFUL and the optimal action. We define the sensitivity ratio at time $t$ as $\alpha_{t}\coloneqq\|x_{t}^{\star}\|_{V_{t}^{-1}}/\|\widetilde{x}_{t}\|_{V_{t}^{-1}}$. We also define the regret ratio at time $t$ as $\mu_{t}\coloneqq\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$.
Recall the Cauchy–Schwarz inequality $\langle{x},{\theta_{1}-\theta_{2}}\rangle\leq\|x\|_{V_{t}^{-1}}\|\theta_{1}-\theta_{2}\|_{V_{t}}$ , intuitively $\alpha_{t}$ serves as a measure of how sensitive the reward $\langle{\widetilde{x}_{t}},{\theta}\rangle$ is compared with that of the optimal action $\langle{x^{\star}_{t}},{\theta}\rangle$, with respect to the change of the parameter $\theta$ in the $V_{t}$-norm.
The intuition behind the definition for $\mu_{t}$ is connecting POFUL and OFUL. It provides an upper bound for the ratio of instantaneous regrets incurred by POFUL and the OFUL-type upper bound. Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$. In this sense, our definition of $\mu_{t}$ and $\alpha_{t}$ can be extended to OFUL. One feasible choice is setting $\mu_{t}\equiv 2$ for OFUL and therefore $\alpha_{t}=(\mu_{t}-1-\iota_{t}-\tau_{t})/(1+\iota_{t}-\tau_{t})=1$ (recall $\iota_{t}=0$ and $\tau_{t}=1$ for OFUL). The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$ and we postpone the proof to Appendix D.3.
@@
A similar notion of (the reciprocal of) our $\alpha_{t}$ can be found in the Discussion section in Abeille et al. (2017), where they indicate their proof of regret bound can be reproduced if the algorithm samples $\widetilde{\theta}_{t}$ that satisfies $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability for some $\alpha>0$. They comment that the set of these parameters doesn’t coincide with the set of optimistic parameters and it remains an open problem whether one could relax the requirement of oversampling the posterior through this observation. Our work answers this question positively in the next section by investigating the property of (the reciprocal of) $\alpha$ using geometric arguments, providing an explanation for LinTS’s empirical success.
@@
We call Theorem 1 an oracle regret bound as $\{\mu_{t}\}_{t\in[T]}$ for general POFUL depends on the unknown system parameter $\theta^{\star}$. In general, they cannot be calculated by the decision-maker. Nevertheless, note that $(1+\iota_{t}-\tau_{t})\geq 0$ for OFUL, LinTS, and Greedy. Suppose we can derive computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$ respectively, then we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well, and Theorem 1 instantly turns into a data-driven regret bound for POFUL. This is indeed our aim in the next section.
5 A Data-Driven Approach to Calibrating POFUL
@@
In this section, we present the main contribution of this work which provides a data-driven approach to calibrating POFUL. We’ll focus on LinTS and Greedy as they lack theoretical results of frequentist regret bound. In these settings, we have $\tau_{t}=0$ and hence $\mu_{t}\coloneqq(1+\alpha_{t})(1+\iota_{t})$. Note that $\iota_{t}$ is the inflation parameter that can be controlled by a decision-maker, the essential point is to find a computable, non-trivial upper bound for the sensitivity ratio $\alpha_{t}$, which turns into an upper bound for the regret ratio $\mu_{t}$ that’s deeply related to the frequentist regret of POFUL.
In the following, we first construct such upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$. Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids. Subsequently, we show how these bounds can be utilized to calibrate POFUL. Specifically, we propose the TS-MR and Greedy-MR algorithms that attain a provable minimax-optimal frequentist regret up to constant factors, and yet keep the majority of their original properties such as computational efficiency and low regret in most instances.
5.1 Upper Bounding the Sensitivity Ratio
@@
We consider two scenarios: (1) Continuous action space and (2) Discrete action space. For sake of a better illustration, we consider $\mathcal{X}_{t}\equiv{\mathcal{S}}_{d-1}$ for the first scenario, which is a standard example of continuous action space, and is the same as the setting considered in Abeille et al. (2017). The analysis of the second scenario is deferred to Appendix A.
Continuous action space.
@@
Note that in this case, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes a closed form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$, we only need to study the property of the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$. A direct idea is to consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the ellipsoids. Specifically, we have
@@
At first glance, one might think that (5) is too conservative to provide useful calibration for POFUL. However, it turns out that the lower bound is pretty tight in practice. As is illustrated in Figure 1. The intuition is that, when POFUL has implemented sufficient exploration so that the confidence ellipsoid $\mathcal{E}_{t}$ is small enough, the potentially optimal actions concentrate in a small cap $\mathcal{C}_{t}$ on the unit hypersphere ${\mathcal{S}}_{d-1}$. All actions within $\mathcal{C}_{t}$ roughly point to the same direction and thus have similar $V_{t}^{-1}$-norm. The following main theorem of this paper derives a upper bound of $\alpha_{t}$ based on this idea. We sketch the proof in Appendix B.
Figure 1: Illustration of the set $\mathcal{C}_{t}$ of potentially optimal actions in $\mathbb{R}^{2}$.
The following assumption simplifies the proof, and we remark that it holds almost surely.
Assumption 4
For all $t$, $V_{t}$ has distinct eigenvalues, i.e. $\lambda_{1}(V_{t})>\lambda_{2}(V_{t})\dots>\lambda_{d}(V_{t})\geq 0$.
@@
Under Assumpsion 4, define $\beta^{PVT}_{t}=\iota_{t}\beta^{RLS}_{t}$ and
@@
When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$. To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small. The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding sensitivity ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1. This provides a plausible explanation for the empirical success of LinTS and Greedy.
@@
5.2 TS-MR and Greedy-MR Algorithms
In the following, we propose the algorithm Linear Thompson Sampling with Maximum (Regret) Ratio (TS-MR). The intuition is to detect the performance of LinTS by estimating $\mu_{t}$ and avoid bad LinTS actions by switching to OFUL occasionally.
Specifically, at each time $t$, TS-MR calculates an estimate $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$. If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action. We remark that setting $\iota_{t}\equiv 0$ yields the corresponding Greedy-MR algorithm. The pseudocode is presented in Appendix E.
By construction, TS-MR and Greedy-MR have $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$. Using Theorem 1, we get the following frequentist regret for them, which is optimal up to a constant factor.
Corollary 1
6 Simulations
(a) example 1
(b) example 2
(c) example 3
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm. At each time $t$, we generate 100 i.i.d. random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$. This is a basic example of standard stochastic linear bandit problems without any extra structure. We set the threshold $\mu=8$ for TS-MR and Greedy-MR. We average simulation results over 20 independent runs. The results are summarized in Figure 1(a). TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$. At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,\mathbb{I}_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0. Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks. This problem is equivalent to a 10-armed contextual bandit. We ser $\mu=12$ for TS-MR and Greedy-MR. Again, we average simulation results over 20 independent runs. The results are summarized in Figure 1(b). In this setting, Greedy performs suboptimally due to a lack of exploration for all arms. Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects bad regret ratio $\widehat{\mu}_{t}$.
Example 3. Prior mean mismatch (Hamidi and Bayati, 2020a).
This is an example in which LinTS is shown to incur linear Bayesian regret. We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},\mathbb{I}_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}},\quad x_{b}=\sum_{i=11}^{3d}\frac{e_{i}}{\sqrt{3d}}-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}}$. It is shown in Hamidi and Bayati (2020a) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$. Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action. We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit. We ser $\mu=12$ for TS-MR and Greedy-MR. We carried out 20 independent runs. The results are summarized in Figure 1(c). We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
7 Conclusion
In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, and Greedy as special cases. Our approach allows for the computation of an instance-dependent frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm. Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest. As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail. We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
Appendix A Discrete Action Space
In this scenario, we consider upper bounding the sensitivity ratio for discrete action space. Compared with the unit hypersphere ${\mathcal{S}}_{d-1}$ considered in Section 5.1 , the main difference is that the oracle optimal action $\theta/\|\theta\|$ for $\theta$ might not lie in the action space $\mathcal{X}_{t}$, so we don’t have a closed form for the attainable optimal action $x^{\star}_{t}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$.
To tackle this problem, we need to characterize the property of all potentially optimal actions. We provide two approaches to solve this problem. This first approach works for the setting where the cardinalities of the action sets are relatively small. The second approach utilizes the statistical properties of the action sets and works for the setting where the cardinalities of the action sets are large.
Small cardinality setting.
Note that the optimistic reward of POFUL when $\tau_{t}=0$ is $\langle{x},{\widetilde{\theta}_{t}}\rangle$, where $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$ is an arbitrary pivot parameter. We have $L_{t}(x)\leq\langle{x},{\widetilde{\theta}_{t}}\rangle\leq U_{t}(x)$, where we define the confidence lower and upper bounds as
$L_{t}(x)=\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}),\quad U_{t}(x)=\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}).$
The potentially optimal $x$ that might optimize $\langle{x},{\widetilde{\theta}_{t}}\rangle$ should satisfy
$U_{t}(x)\geq\max_{y\in\mathcal{X}_{t}}\ L_{t}(y),$
i.e. $x$ is not dominated by any other action in $\mathcal{X}_{t}$. Therefore, define
$\mathcal{C}_{t}(\beta)\coloneqq\{x\in\mathcal{X}_{t}:\langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta\geq\max_{y\in\mathcal{X}_{t}}\ \langle{y},{\widehat{\theta}_{t}}\rangle-\|y\|_{V_{t}^{-1}}\beta\}.$
An upper bound for $\alpha_{t}$ can be calculated as
$\alpha_{t}\leq\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t}^{RLS}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}.$ (6)
Large cardinality setting.
In this setting, we characterize the distance between $\theta/\|\theta\|$ and $x^{\star}_{t}(\theta)$. We first introduce the notion of Euclidean covering radius.
Definition 6
The Euclidean covering radius $\rho_{t}$ of a set $\mathcal{X}_{t}\subseteq{\mathcal{S}}_{d-1}$ is given by
$\rho_{t}=\rho(\mathcal{X}_{t},{\mathcal{S}}_{d-1})\coloneqq\max_{y\in{\mathcal{S}}_{d-1}}\min_{x\in\mathcal{X}_{t}}\|y-x\|.$
With the help of the Euclidean covering radius, we prove the following variant of Theorem 2. The proof is postponed to Appendix D
Theorem 3
Under Assumpsion 4, let $\{\rho_{t}\}$ be the sequence of Euclidean covering radius of $\mathcal{X}_{t}$, define $\Psi_{t},\Phi_{t}$ as in Theorem 2. Then for all $t\in[1,T]$, conditioned on event $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$ , it holds that
$\alpha_{s}\leq\frac{\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})}{\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})}$ (7)
for all $s\leq t$.
To justify this new lower bound, it remains to show that $\rho_{t}$ is small enough under proper conditions. For the sake of a better illustration, we focus on action sets that consist of uniformly and independently distributed points on ${\mathcal{S}}_{d-1}$. The following proposition shows that the Euclidean covering radius has an high-probability upper bound that decays at the order of $\widetilde{\mathcal{O}}(N^{-1/d+o(1)})$, where $N$ is the cardinality of actions.
Proposition 6 (Corollary 2.9 in Reznikov and Saff [2016])
Suppose $\mathcal{X}$ is a set of $N$ uniformly and independently distributed random points on the unit sphere ${\mathcal{S}}_{d-1}$, for any $\kappa>0$, there exist constants $C_{\kappa,d}>0$, such that
$\mathbb{P}\left[\rho(\mathcal{X}_{t},{\mathcal{S}}_{d-1})\leq C_{\kappa,d}\left(\frac{\log N}{N}\right)^{1/d}\right]\geq 1-N^{-\kappa}.$
Appendix B Sketch of the Proof of Theorem 2

### v3

@@ 4.1 An Instance-Dependent Regret Bound for POFUL
In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$. The following proposition bounds the instantaneous regret of POFUL and we defer the proof to Appendix 10.2.
@@
On the right-hand side, since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\left\|\widetilde{x}_{t}\right\|_{V_{t}^{-1}}^{2}$ to get an upperbound of the regret. To fix the problem due to this discrepancy, the key point is connecting $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ in terms of the $V_{t}^{-1}$-norm. This motivates the following definition.
@@
For each $t\geq 1$, let $\widetilde{x}_{t}$ and $x_{t}^{\star}$ respectively denote the action chosen by POFUL and the optimal action. We define the uncertainty ratio at time $t$ as $\alpha_{t}\coloneqq\|x_{t}^{\star}\|_{V_{t}^{-1}}/\|\widetilde{x}_{t}\|_{V_{t}^{-1}}$. We also define the (instantaneous) regret proxy at time $t$ as $\mu_{t}\coloneqq\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$.
Note that $\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}$ holds with high probability, we have that $\|x\|_{V_{t}^{-1}}$ essentially determines the length of the confidence interval of the reward $\langle{x},{\theta^{\star}}\rangle$. Hence, $\alpha_{t}$ serves as the ratio of uncertainty degrees of the reward obtained by the optimal action $x^{\star}_{t}$ and the chosen action $\widetilde{x}_{t}$.
The intuition behind the definition for $\mu_{t}$ is constructing a regret upper bound similar to that of OFUL. Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$. In this sense, $\mu_{t}$ is a proxy of the instantaneous regret incurred by POFUL at time $t$. Moreover, OFUL can be regarded as a POFUL algorithm whose $\mu_{t}$ is fixed at 2, and we could extend the definition of $\alpha_{t}$ to OFUL by solving $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$ and set $\alpha_{t}\equiv 1$ for OFUL (recall $\iota_{t}\equiv 0$ and $\tau_{t}\equiv 1$ for OFUL).
The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$ and we postpone the proof to Appendix 10.3. It provides an oracle but general frequentist regret upper bound for all POFUL algorithms.
@@
We call Theorem 1 an oracle regret bound as $\{\mu_{t}\}_{t\in[T]}$ for general POFUL depends on the unknown system parameter $\theta^{\star}$. In general, they cannot be calculated by the decision-maker. Nevertheless, note that $(1+\iota_{t}-\tau_{t})\geq 0$ for OFUL, LinTS, and Greedy. Suppose we can derive computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$ respectively, then we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well, and Theorem 1 instantly turns into a data-driven regret bound for POFUL, and could be utilized later for the course correction. This is indeed our aim in the next section and we achieve it using geometric arguments.
@@
A similar notion of (the reciprocal of) our $\alpha_{t}$ can be found in the Discussion section in Abeille et al. (2017), where they indicate their proof of regret bound for LinTS can be reproduced if LinTS samples $\widetilde{\theta}_{t}$ satisfying $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability for some $\alpha>0$. They comment that the set of these parameters doesn’t coincide with the set of optimistic parameters and it remains an open problem whether one could relax the requirement of oversampling the posterior through this observation. Our work answers this question positively in the next section by investigating the property of (the reciprocal of their) $\alpha$ using geometric arguments, providing an explanation for LinTS’s empirical success.
5 A Data-Driven Approach
@@
In this section, we present the main contribution of this work which provides a data-driven approach to calibrating POFUL. Note that $\iota_{t}$ and $\tau_{t}$ are parameters of POFUL that can be controlled by a decision-maker, the essential point is to find a computable, non-trivial upper bound $\widehat{\alpha}_{t}$ for the uncertainty ratio $\alpha_{t}$, which turns into an upper bound $\widehat{\mu}_{t}$ for the regret proxy $\mu_{t}$ that’s deeply related to the frequentist regret of POFUL.
Typically, we can check $\widehat{\alpha}_{t}$ is no less than one because we might have $\widetilde{x}_{t}=x_{t}^{\star}$. As a result, it holds that $\widehat{\mu}_{t}=(1+\widehat{\alpha}_{t})(1+\iota_{t})+(1-\widehat{\alpha}_{t})\tau_{t}\geq(1+\widehat{\alpha}_{t})(1+\iota_{t})$, i.e., non-zero $\tau_{t}$ gives a larger regret proxy. Given this observation, we’ll focus on scenarios where $\tau_{t}\equiv 0$. This avoids the large regret proxies and computational inefficiencies seen in OFUL when determining the action $\widetilde{x}_{t}$. Such scenarios include LinTS and its variants like TS-Freq as well as Greedy - all of which are standard algorithms that still lack theoretical regret guarantee results.
In the following, we construct such upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for both discrete and continuous action scenarios.
5.1 Discrete Action Space
We start with the discrete action space scenario to illustrate our method. It turns out that upper bounding $\alpha_{t}$ is done by simply comparing the confidence interval lengths of all potentially optimal actions.
To see this, recall for an arbitrary action $x\in\mathcal{X}_{t}$, by Proposition 1, with high probability its expected reward is bounded as
$L_{t}(x)\coloneqq\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime})\leq\langle{x},{{\theta}^{\star}}\rangle\leq\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime})\eqqcolon U_{t}(x).$
The confidence interval length is proportional to $\|x\|_{V_{t}^{-1}}$. Therefore, to characterize the range of $\|x_{t}^{\star}\|_{V_{t}^{-1}}$, we only need to identify the set of potentially optimal actions and calculate the minimal $V_{t}^{-1}$-norm among all actions in that set.
Note that for an action $x\in\mathcal{X}_{t}$ to remain potentially optimal, it cannot be dominated by another action. I.e., we require $U_{t}(x)\geq\max_{y\in\mathcal{X}_{t}}L_{t}(y)$. Therefore, define
$\mathcal{C}_{t}(\beta)\coloneqq\{x\in\mathcal{X}_{t}:\langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta\geq\max_{y\in\mathcal{X}_{t}}\ \langle{y},{\widehat{\theta}_{t}}\rangle-\|y\|_{V_{t}^{-1}}\beta\},$
the set of potentially optimal actions is given by $C_{t}(\beta_{t}^{RLS}(\delta^{\prime}))$. Similarly, the set of actions that might be chosen by POFUL is given by $\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))$, where we let $\beta^{PVT}_{t}\coloneqq\iota_{t}\beta^{RLS}_{t}$. The inflation parameter $\iota_{t}$ turns out to control the conservativeness when eliminating actions.
By comparing the range of $V_{t}^{-1}$-norm for both sets, an upper bound $\widehat{\alpha}_{t}$ for the uncertainty ratio $\alpha_{t}$ is given by
$\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t}^{RLS}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}.$ (5)
5.2 Continuous action space.
When it comes to the generic settings where the number of actions is large or when the set of actions is a continuous set in $\mathbb{R}^{d}$, the upper bound above is inefficient or even infeasible to compute. In this section, we tackle this issue. Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids, providing upper bounds that can be computed efficiently.
@@ 5.1 Discrete Action Space
For the sake of a better illustration, we consider $\mathcal{X}_{t}\equiv{\mathcal{S}}_{d-1}$ for this scenario, which is a standard example of continuous action space, and is the same as the setting considered in Abeille et al. (2017). We remark that for this specific setting, the problem is still hard. This is because we don’t have a closed-form solution for the set of potentially optimal actions.
@@
In this setting, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes a closed form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$. To upper bound $\alpha_{t}$, a direct idea is to consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$. Specifically, we have
@@
As is illustrated in Figure 1. The set of potentially optimal actions $\mathcal{C}_{t}$ is the projection of the confidence ellipsoid $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$. It’s hard to get a closed-form expression for $\mathcal{C}_{t}$, so we cannot directly calculate the range of $V_{t}^{-1}$-norm of actions in $\mathcal{C}_{t}$. Nevertheless, we can approximate the range by investigating the geometric properties of the ellipsoids.
The intuition here is that, when POFUL has implemented sufficient exploration so that $\mathcal{E}_{t}$ is small enough, $\mathcal{C}_{t}$ concentrates accordingly to a small cap on ${\mathcal{S}}_{d-1}$. All actions within $\mathcal{C}_{t}$ point to similar directions and thus have similar $V_{t}$-norm. (Note that for a unit vector $x$, we have that $\|x\|_{V_{t}}$ can be written as a weighted summation of the eigenvalues of $V_{t}$, where the weights are determined by the direction of $x$.) Hence we can bound the range of $V_{t}$-norm by investigating the concentration of $\mathcal{C}_{t}$, which is pure geometry, and we’ll use the range of $V_{t}$-norm to determine the range of $V_{t}^{-1}$-norm.
Figure 1: Illustration of the set $\mathcal{C}_{t}$ of potentially optimal actions in $\mathbb{R}^{2}$. (a): $\mathcal{C}_{t}$ is the projection of $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$. (b): As the decision-maker collects more and more data, the confidence ellipsoid $\mathcal{E}_{t}$ shrinks accordingly (here different color corresponds to different levels of exploration). The potentially optimal actions point to similar directions, which essentially determine their $V_{t}$-norm. Hence it’s plausible one could get a proper range of their $V_{t}$-norm using geometry.
The following main theorem of this paper derives an upper bound of $\alpha_{t}$ based on this idea. We sketch the proof in Section 5.3.
@@
Suppose $\mathcal{X}_{t}={\mathcal{S}}_{d-1}$ for all $t\in[T]$. Define $\beta^{PVT}_{t}\coloneqq\iota_{t}\beta^{RLS}_{t}$ and
@@
When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$ (empirically validated in Appendix 11). To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small. The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding uncertainty ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1. This provides a plausible explanation for the empirical success of LinTS and Greedy.
@@
Remark 3
Clearly, one can further improve the bound in Equation (6) by plugging in the selected action $\widetilde{x}_{t}$. This is true for standard algorithms such as OFUL, LinTS, and Greedy. Explicitly, the uncertainty ratio can be bounded using $\widetilde{x}_{t}$ and Theorem 2 as
$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\|\widetilde{x}_{t}\|_{V_{t}^{-1}}}\leq\frac{\Phi_{t}}{\|\widetilde{x}_{t}\|_{V_{t}^{-1}}}.$ (7)
However, for the variants of POFUL that we are going to propose in the Section 6, it is necessary to calculate the bound before selecting the action $\widetilde{x}_{t}$. Therefore, the bound shouldn’t depend on $\widetilde{x}_{t}$.
5.3 Sketch of the Proof